OpenStudy (anonymous):

please help ill give medals & fans Use a both elimination and substitution to solve this system of linear equations. Type your answer as an ordered triplet, like this: (-2, 4, 13) 2x−10y+1=z 4x + y = z + 18 -2x - 8z = y - 1

OpenStudy (anonymous):

can u make sure you types your equations right

OpenStudy (anonymous):

@lacrosseplayer22 Use a both elimination and substitution to solve this system of linear equations. Type your answer as an ordered triplet, like this: (-2, 4, 13) 2x−10y+1=z 4x + y = z + 18 -2x - 8z = y - 1

OpenStudy (anonymous):

give me a min.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x=4 I think

OpenStudy (anonymous):

@Venny is x=4?

OpenStudy (anonymous):

y=1 if x=4

OpenStudy (anonymous):

z=-1 if x=4 and y=1

OpenStudy (anonymous):

plz correct me if im wrong

OpenStudy (anonymous):

I would have to show a lot of work but basically I substituted the variables into each equation......

OpenStudy (anonymous):

sorry im not very good at explaining.

OpenStudy (anonymous):

am I right?

OpenStudy (anonymous):

@Lovee_Stasiaa do you have multiple choice options?

OpenStudy (anonymous):

@lacrosseplayer22 sorru i was feeding my daughter you were right (: